# Radial Gradients and Intrinsic Scatter in MaNGA Galaxies

### v1

@@ 4.2.1 Age gradients
In Fig. 6 we showed how the age is changing with $\sigma$. In this section, we concentrate on how radial gradients in age are changing for galaxies within a particular $\sigma$ bin ($\sigma$ bins are described in Table 2 or 4). In Fig. 13 we plot the mean age of each $\sigma$ bin against the log(R/Re) values. The errorbars for each bin comes from the mean of all the errors within that bin. The errorbars are larger in case of age. On average, we find a flat or very small negative radial gradient in age for all $\sigma$ bins. However, the scatter is large, implying that while many galaxies are younger in the outskirts, nearly as many are younger in the centers. Only the highest $\sigma$ bin, with the least-secure counting statistics, shows a situation where the clear majority of the galaxies have older-looking centers and younger outskirts.
The top panel of Fig. 14 shows the radial gradient values for individual galaxies plotted against $\sigma$. The gradient values are predominantly centered near zero. In the bottom panel of Fig. 14 we plotted the mean radial gradient value for each $\sigma$ bin against the middle value of corresponding $\sigma$ bin. The least $\sigma$ bin has near zero mean radial gradient whereas the other $\sigma$ bins have a near constant value of -0.05 dex/log(R/Re). It should be noted that the gradient error for each bin is not the standard deviation of each bin, but the mean of the radial gradient errors for that bin, which is meant to indicate a typical one-galaxy error.
We need to mention a couple of subtle points as regards Figures 13 and 14. While calculating the mean values in Fig. 13 and bottom panel of Fig. 14, we have omitted outliers using the robust Z-score method. Data points with a Z-score value greater than 3.0 we considered outliers. Small differences arise when comparing the gradients for all the $\sigma$ bins from Fig. 13 and the radial gradient values shown in the bottom panel of Fig. 14. This is due to methodology. In Fig. 13, we calculate the gradient after we calculate mean age (after removing outliers) for each $\sigma$ bin but in Fig. 14 we first calculate the radial gradient for all galaxies and then calculate the mean. It is generally expected that gradient of the mean will not equal the mean of the gradient. Also, while removing outliers, the methods remove different data points and thus the figures do not produce identical results. These comments are true for Fig. 15 & Fig. 16 pair, and Fig. 17 & figures 18, 19, 20, and 21 pairs.
@@
Fig. 15 shows how the mean [Fe/H] for different $\sigma$ bins (color coded) is changing with radius with log(R/Re) as the proxy for radius. Qualitatively, a strong negative radial gradient is observed for all $\sigma$ bins in the sense that very few galaxies show a positive gradient at any bin in $\sigma$. Fig. 16 shows the radial gradient values in dex per decade. Individual radial gradients are plotted against corresponding $\sigma$ values on the top panel of Fig. 16. Similar to §$\S$4.2.1, the galaxies are grouped in five $\sigma$ bins and mean gradient values are plotted against $\sigma$ in the bottom panel of Fig. 16. The errors are again the mean of individual errors within each $\sigma$ bin.
@@
Compared to §$\S$4.2.1, the radial gradients in [Fe/H] have higher negative values and smaller intrinsic scatter (we will shortly show that most of the scatter is astrophysical, not observational in origin). From Fig. 16, the highest $\sigma$ bin shows a gradient value of -0.13 dex/log(R/Re) which implies that the [Fe/H] drops by 0.13 dex when we move from the innermost annulus to the 5th annulus. The fact that the smallest galaxies, which may or may not be merger remnants, have the shallowest [Fe/H] gradients while the largest galaxies, which are definitely shaped primarily by mergers, have the strongest gradients is a point for discussion.
@@ 4.2.3 C, N, Na, and Mg gradients
Fig. 17 shows how light element abundances are changing with respect to Fe, radially, in five different $\sigma$ bins. The errorbars associated with each bin is the mean of errors in radial gradient values for the galaxies within that bin. The figure shows once again that the abundances increase with increasing $\sigma$ value (see also Fig. 8) at all radial values. Na appears to have the highest gradients for all $\sigma$ values, though due to axis scaling this might most easily be seen in Table 4. Another point worth mentioning here is that not just the abundances, but also the magnitude of the radial gradient steepens with increasing $\sigma$ value. This implies that more massive galaxies tend to have steeper negative radial gradient in abundances, especially for Na.
@@
Figures 18, 19, 20, and 21 show how the individual radial gradient (top panel) of each galaxy and the mean of radial gradients for each $\sigma$ bin (bottom panel) changes with $\sigma$ for C, N, Na, and Mg respectively. The mean values for C, N, and Mg are close to zero for all the $\sigma$ bins with decreasing trend with $\sigma$ (the decreasing trend is very prominent for Na and Mg). For Na, there is a strong dependence of the radial gradient value with $\sigma$. Not only there is a strong gradient in radial gradient but the absolute values of the radial gradients are also high at higher $\sigma$ values for Na (Fig. 20). For the highest $\sigma$ bin in Fig. 20, a negative gradient value of 0.15 signifies that if we move to the 5th annulus from the center of the galaxy, there is a drop in Na abundance by 0.15 dex. For Mg, although the higher $\sigma$ bins have a more negative mean radial gradient value, the absolute values are on the lower side. The low-$\sigma$ galaxies have near-zero gradients in individual light element abundances , though most of them have negative [Fe/H] gradients.
@@
Since we are treating each galaxy as an individual entity in this work, we wanted to look into how the radial gradient values in these four elements correlate with each other (if there is any correlation at all). In order to check any such correlation, we plotted the individual radial gradients in different elements with each other and calculated the Spearman correlation. A positive Spearman coefficient signifies a positive correlation with values closer to one meaning very high positive correlation, whereas a negative value indicates a anti-correlation with values close to negative one signifying the most anti-correlation.
Fig. 22 shows the plots of radial gradients in C, N, Na, and Mg plotted against each other for all possible combinations of these four elements. The radial gradients in C and Na seem to correlate the most ($\rho$=0.36) compared to others. But a low $\rho$ value signifies little positive correlation. Mg has a very similar correlation ($\rho\sim$0.3) with all the rest of the elements (C, N, and Na). N and C are the least correlated pair among all of them with a $\rho$ value of 0.13, this is almost equivalent to having no correlation (signified by a $\rho$ value near zero). Na and N have the second least correlation value at $\rho$=0.21.
Figure 22: The radial gradient values in C, N, Na, and Mg for each individual galaxy are plotted against one another. The legend shows the Spearman coefficient value for each combination of elements. The errorbars are added for each data point as well.
@@ 5 Discussion
In this work, we derived the stellar population parameters (age, [Fe/H], [C/R], [N/R], [Na/R], and [Mg/R]) and their radial gradients for 2417 individual MaNGA galaxies at different radial distances from the center of the galaxy. In §$\S$4.1, we show how different stellar population parameters are changing with $\sigma$ and [Fe/H] whereas in §$\S$4.2, we illustrate the dependence of radial gradients on $\sigma$. In this section we compare some of our key results with previous work and also provide a broader implication of this work in terms of galaxy formation and evolution.
@@ 5.1 Stellar Population Parameters
Fig. 6 shows how the individual age of galaxy (for the 3rd annulus; bottom panel) and mean age (within each $\sigma$ bin; top panel) changes with $\sigma$. We see a positive correlation between the two with the age steadily increasing with the size of the galaxy. This is also evident from the positive slope of 0.59 in the bottom panel of Fig. 6. This slope is in good agreement with Parikh et al. (2019) (their Fig. 7) and Worthey et al. (2014) (their Fig. 3). Although Worthey et al. (2014) does not provide with a trendline fit to the data, the extent of the age at two ends of $\sigma$ is similar to that of our result. Conroy et al. (2014)’s Fig. 18 shows variation of different parameters with $\sigma$ with a gradient of around 0.55 for log(age) vs log($\sigma$) relationship. This difference in the gradient (compared to our 0.59) rises probably because of difference in stellar population models used for these two works. Johansson et al. (2012) reports a gradient value of 0.73 for their log(age) vs log($\sigma$) relationship for around 4000 SDSS ETGs. Within the scope of errors, we also reproduced similar results as Johansson et al. (2012) with our individual MaNGA galaxies. These various good agreements are interesting because of the volatile oxygen-age degeneracy Worthey et al. (2022) along with our assumption that [Mg/R] = [O/R]; with such a large increase in model flexibility, there was no a priori reason that the age scale would remain unchanged. This trend of larger galaxies having older age is valid irrespective of the density of environment according to Thomas et al. (2005). These considerations support an anti-hierarchical picture where larger (more massive) ETGs were formed at an earlier stage of the evolution of the Universe.
Shifting our focus to [Fe/H], Fig. 7 showed how [Fe/H] varies as a function of $\sigma$ for both individual galaxies (3rd annulus; bottom panel) and $\sigma$ bins (top panel). In conjunction with Fig. 6, we can also conclude that older galaxies tend to have slightly lower [Fe/H]. We emphasize again that due to innovations in our modeling (isochrones that shift in temperature for individual elements and metallicity-composite base models) our results should not necessarily track previously published results. That said, a similar trend is also found in Trager et al. (2000b) where for a given $\sigma$, older galaxies are found to have low [Fe/H]. Although the opposite trend is shown in some papers (Conroy et al., 2014), many studies show a similar flat to negative dependence on [Fe/H] when plotted against $\sigma$ (Johansson et al., 2012). This behavior can be attributed to several factors.
@@
In Fig. 8 we see a trend of increasing abundance for all elements explored in this work (C, N, Na, and Mg). Previous work on elemental abundances in ETGs shows a very similar trend (Parikh et al., 2021; Johansson et al., 2012; Conroy et al., 2014; Graves et al., 2007). Apart from Mg, none of the elements explored in this work can be termed as $\alpha$ elements. Each of them have different pathways of formation.
@@
Nitrogen is expelled from supernova explosions, but is mostly formed through the CNO cycle in stars from carbon. Some carbon is also produced by supernovae, but most comes from the triple alpha process. For both elements, stellar winds during the late stages of evolution enrich the nearby interstellar medium. For stellar populations, carbon production is dominant in stars with lower mass (1-3 M) whereas for nitrogen we need higher mass stars (4-8 M) (Chiappini et al., 2003). In Milky Way stars, [C/Fe] hovers near zero as a function of [Fe/H], while [N/Fe] rises with metallicity (Jönsson et al., 2018). In this work, we see that N is under-abundant (on average by 0.25 dex) compared to C, and compared always to the solar ratio. Conroy et al. (2014) reported similar abundance values and gradient for C and N.
@@
Not included in the models explicity is the effect of first dredge-up on the first-ascent red giant branch. Nitrogen abundance increases by $\sim 0.4$ dex relative to C on the upper RGB [see, e.g., Fig 2 of Martig et al. (2016)]. If this effect were included, nitrogen’s spectral signatures would strengthen and the inferred N abundance would drop even lower in Fig. 12, putting [N/C] and [N/Fe] significantly low compared to the Milky Way. The effect is included implicitly, in that Milky Way stars were used to empirically zero the spectral library, and this effect is built in, masking as a surface gravity effect. That said, and keeping also in mind the N-rich globular cluster system in M31 that contrasts so strongly with our own cluster system (Worthey, 1998) and the small dynamic range of current galaxy formation model predictions (Fig. 31), it is safe to say that N is telling us much about chemical evolution that we are not yet equipped to interpret.
[…]
@@
A remarkable finding from this work is shown in Fig. 9. The effect shows more clearly because we did not stack spectra. What we see is that there is a very prominent increase in slope of chemical abundances when plotted against $\sigma$ for log($\sigma$)$>$2.0. This slope change seems not to have been mentioned in previous literature on this subject. Possibly related, however, is that Gallazzi et al. (2005) found a transition from low mass to massive galaxies in the mass-metallicity relation at around log($\sigma$)=2.3. Evidently, low mass galaxies are less efficient in producing light metals and retaining them compared to massive galaxies. Different factors could lead to this, including star formation efficiency, galactic winds, and the IMF.
[…]
5.2 Radial Gradients in Stellar Population Parameters
@@
The radial gradients in different stellar population parameters provide insight into the formation and evolution of galaxies. In this work, in addition to looking at the trends of stellar population parameters with $\sigma$, we also investigate their radial gradients and their trends with $\sigma$.
@@
In Fig. 13 we show how the mean age for each $\sigma$ bin changes radially from the center of the galaxy to 0.7 Re. We find a somewhat shallow decreasing trend of mean age with radius on average, but the scatter is much larger than the mean trend. This result is also validated through Fig. 14. There are studies that report nearly zero radial gradient in age (Mehlert et al., 2003; Sánchez-Blázquez et al., 2007; Kuntschner et al., 2010). Some other studies show negative mean radial gradient in age (Rawle et al., 2008, 2010; Parikh et al., 2021). Our result, showing a modest radial gradient in age, is consistent with these results within the scope of error. Most of the previous studies on radial gradient in age consider ages out to R=Re, but in this work we calculated spatial information to R=0.7 Re. A lack of age gradient might imply that stellar populations in ETGs were formed in bursts that spanned large swaths of galactocentric radius. In other words, star formation tends to be global, not outside-in or inside-out. However, it needs to be remembered that our results apply to the average. Few galaxies conform to the average. Slightly more than half are younger in the outskirts (0.7 Re for this study) and slightly less than half are younger in the center. Only a minority of galactic accretion events drive gas to galaxy centers, it seems.
@@
We find radial gradients in [Fe/H] $\approx-0.10$ [dex per decade, [Fe/H]/log(R/Re)] for the middle $\sigma$ bin with substantial dependence on $\sigma$ (see Fig. 16). This value is very similar to values reported by Parikh et al. (2021) and Mehlert et al. (2003) but of shallower slope than some other works (Kobayashi, 2004; Kuntschner et al., 2010). Earlier estimates averaged a slope of $-0.2$ dex per decade for medium and large ETGS (Henry & Worthey, 1999). We see a considerable scatter around zero (see top panel of Fig. 16) with values as negative as -0.5 [Fe/H]/log(R/Re) and as positive as 0.25 [Fe/H]/log(R/Re) (not considering outliers). Our trend of radial [Fe/H] gradient with $\sigma$ qualitatively matches with Spolaor et al. (2009) at low $\sigma$ values (log($\sigma$)$<$2.0) and with Bekki & Shioya (1999) for higher $\sigma$ values. A negative [Fe/H] gradient favors an ‘inside-out’ mode of chemical evolution in ETGs if star formation and enrichment occur on a timescale similar to the accretion itself. This conclusion is similar to that of Goddard et al. (2017) where they find similar results with 505 ETGs. The [Fe/H] radial gradient reported here could also be explained as a consequence of hierarchical merging events. If they involve gas, mergers can trigger intense central starbursts, which significantly enrich the core regions with metals leading to a metallicity distribution where the central areas exhibit higher metal content than the outer regions (Kobayashi, 2004). If not subsumed into star formation right away, gas loses energy and angular momentum, funneling into the central regions where concentrated star formation enriches the core with metals more significantly than the outskirts (Montes et al., 2014). The gradual increase of [Fe/H] radial gradient with $\sigma$ can also be explained by the amplification of these processes for larger galaxies. The monolithic collapse hypothesis on the other hand predicts a much higher [Fe/H] gradient (Kawata, 2001; Pipino et al., 2010) than what is reported here. If monolithic collapses are a common formation channel, then we also require major “dry” merger events to explain the much flatter slope in [Fe/H].
@@
Using a linear scale for R/Re, the negative radial gradient in age is 0.04 dex per R/Re for the lowest $\sigma$ bin and remains constant at an increased value of 0.11 dex per R/Re for the rest of the $\sigma$ bins. We quantify the goodness of fit for both logarithmic and linear cases using the coefficient of determination aka the R2 value. The R2 values for each $\sigma$ bin for the linear case are (in order of increasing $\sigma$): 0.52, 0.93, 0.97, 0.94, and 0.95. On the other hand, for the logarithmic case, the R2 values are: 0.26, 0.70, 0.77, 0.71, 0.85. Thus, in general, the linear case shows better fit. Similar analyses were also performed for the radial gradients in [Fe/H]. The negative radial gradient in [Fe/H] (using linear scale for R/Re) increases from 0.12 dex per R/Re to 0.23 dex per R/Re with R2 values of 0.96, 0.97, 0.97, 0.98, and 0.98 for different $\sigma$ bins (in order of increasing $\sigma$). Using the logarithmic scale the R2 values are 0.79, 0.81, 0.91, 0.93, and 0.84, again showing better fit for the linear case. Thus for both age and [Fe/H] we find that a linear physical scale in kpc connects to the spatial dependencies of chemical evolution better than a logarithmic one. For example, this might mean that power-law dependencies such as depth of potential well or dynamical crossing time are less important than turbulence or other processes with a relatively fixed sphere of influence.
@@
Fig. 17 shows the radial trend of all of the elements explored in this work. Below we’ll discuss each of the elements individually.

### v2

@@ 4.2.1 Age gradients
This section focuses on the radial gradients in age across galaxies within specific $\sigma$ bins, as detailed in Table 4. In Fig. 14, the mean age for each $\sigma$ bin is plotted against log R/Re (color-coded by $\sigma$ bins), with error bars representing the mean uncertainty within each bin. On average, radial age gradients are flat or slightly negative ($\sim$-0.04 dex per log(R/Re) across all $\sigma$ bins, though significant scatter indicates that many galaxies have younger outskirts, while a comparable number are younger at their centers. Only the highest $\sigma$ bin, with limited statistical reliability, consistently shows older centers and younger outskirts.
The top panel of Fig. 15 shows radial age gradient values for individual galaxies as a function of $\sigma$, with most values clustering near zero. The bottom panel plots the mean radial gradient for each $\sigma$ bin against the bin’s central $\sigma$ value. The lowest $\sigma$ bin exhibits a near-zero mean radial gradient, while other bins maintain a near-constant gradient of -0.04 dex/log(R/Re). Notably, the gradient error for each bin is derived from the mean of individual radial gradient uncertainties within the bin, reflecting a typical single-galaxy error, rather than the standard deviation of the bin.
Several methodological nuances warrant clarification regarding Figures 14 and 15. Mean values in Fig. 14 and the bottom panel of Fig. 15 exclude outliers identified using the robust Z-score method, with data points exceeding a Z-score of 3.0 classified as outliers. Discrepancies between the gradients in Fig. 14 and the mean radial gradients in the bottom panel of Fig. 15 arise from different computational approaches. In Fig. 14, the gradient is computed after calculating the mean age (post-outlier removal) for each $\sigma$ bin, whereas in Fig. 15, radial gradients are first calculated for individual galaxies before averaging. As the gradient of the mean does not generally equal the mean of the gradients, and different data points are excluded as outliers, the results are not identical. These considerations also apply to the pairs Fig. 16 and Fig. 17, as well as Fig. 18 and Figures 19, 20, 21, and 22.
@@
Fig. 16 shows the variation of mean [Fe/H] across different $\sigma$ bins (color-coded) as a function of radius, using log(R/Re) as the radial proxy. At a glance, a pronounced negative radial gradient is consistently observed across all $\sigma$ bins. Fig. 17 shows the radial gradient values in dex per decade. The top panel of Fig. 17 plots individual radial gradients against their corresponding $\sigma$ values. Consistent with $\S$4.2.1, galaxies are grouped into five $\sigma$ bins, and the mean gradient values for each bin are plotted against $\sigma$ in the bottom panel of Fig. 17. The error bars represent the mean of individual uncertainties within each $\sigma$ bin.
@@
Compared to $\S$4.2.1, the radial gradients in [Fe/H] have higher negative values and smaller intrinsic scatter (we will show in $\S$5.3 that most of the scatter is astrophysical, not observational in origin). In Fig. 17 the highest $\sigma$ bin shows a gradient value of -0.16 dex/log(R/Re) which implies that the [Fe/H] drops by 0.16 dex when we move from the innermost annulus to the 5th annulus. The fact that the smallest galaxies, which may or may not be merger remnants, have the shallowest [Fe/H] gradients while the largest galaxies, which are definitely shaped primarily by mergers, have the strongest gradients is a point for discussion (see $\S$5.2).
@@ 4.2.3 C, N, Na, and Mg gradients
Fig. 18 shows how [X/R] is changing radially, in five different $\sigma$ bins. The errorbars associated with each bin is the mean of errors in radial gradient values for the galaxies within that bin. The figure shows once again that the abundances increase with increasing $\sigma$ value (see also Fig. 10) at all radial values. Na appears to have the highest gradients for all $\sigma$ values, though due to axis scaling this might most easily be seen in Table 4. Another point worth mentioning here is that apart from Na, all the other light elements considered in this work show $\sim$-0.03 dex/log R/Re radial gradient value across all the $\sigma$ bins (very similar to the radial gradient in age).
@@
Figures 19, 20, 21, and 22 shows the variation of radial gradients for C, N, Na, and Mg, respectively, with respect to $\sigma$. The top panels illustrate individual galaxy radial gradients, while the bottom panels present the mean radial gradients for each $\sigma$ bin. For C, N, and Mg, the mean radial gradients are approximately zero ($\sim$-0.03) across all $\sigma$ bins, showing minimal dependence on $\sigma$. In contrast, Na exhibits a strong dependence of radial gradient on $\sigma$, with both a pronounced variation in gradient and higher absolute gradient values at high $\sigma$ values (Fig. 21). Specifically, in the highest $\sigma$ bin of Figure 21, a mean negative gradient of 0.15 dex/log R/Re indicates a decrease in Na abundance by 0.15 dex from the galaxy center to the 5th annulus. Low-$\sigma$ galaxies exhibit near-zero gradients for individual light element abundances, despite predominantly negative [Fe/H] gradients.
@@ 5 Discussion
In this work, we derived stellar population parameters (mean light-weighted age, ADF peak [Fe/H], [C/R], [N/R], [Na/R], and [Mg/R]) and their radial gradients for 2968 individual MaNGA galaxies. Here, we review and amplify key results for (1) stellar population parameter trends, (2) gradient trends, and (3) astrophysical scatter. Additionally, in $\S$5.4 we compare our results with TNG, a galaxy evolution simulation. We compare key results with previous work and mention some broader implications for future focus.
@@ 5.1 Stellar Population Parameters
An ingredient currently missing from the models is blue straggler stars. Inclusion of these bluer stars will increase the inferred ages, likely by Gyr (Trager et al., 2000a). The ages as reported here are by no means the last word on the subject.
Shifting our focus to [Fe/H], Fig. 9 showed how [Fe/H] varies as a function of $\sigma$ for both individual galaxies (3rd annulus; bottom panel) and $\sigma$ bins (center of rectangle; top panel). In conjunction with Fig. 8, we can also conclude that older galaxies tend to have slightly lower [Fe/H]. We emphasize again that due to innovations in our modeling (isochrones that shift in temperature for individual elements and metallicity-composite base models) our results should not necessarily track previously published results. That said, Trager et al. (2000b) echoes our finding that for a given $\sigma$, older galaxies are found to have low [Fe/H]. In the bottom panel of Fig. 23, we compare our log $\sigma$ - [Fe/H] relation with that of Johansson et al. (2012) (orange line) and Conroy et al. (2014) (blue line). Neither of these works report a ‘kink’ in this relation. Although Johansson et al. (2012) show a flat to negative dependence of [Fe/H] with $\sigma$, the opposite trend is shown by Conroy et al. (2014).
The behavior of decreasing (or flat) [Fe/H] with $\sigma$ can be accommodated within current galaxy evolution theory. A possible factor that might reduce [Fe/H] in galaxies is a high star formation rate (SFR) of the galaxy (Lara-López et al., 2010) followed by swift quenching before Type Ia products contribute. Davé (2008) have shown that galaxies with higher stellar mass tend to have higher SFR within a redshift ($z$) range of 0-2. The higher SFR would logically produce stronger outflows within a galaxy that further reduces its [Fe/H] (Cicone et al., 2016), although a deeper gravitational potential well would work against that. If star formation occurs in stronger bursts in more massive (and thus older) ETGs, then massive ETGs formed stars early in the Universe and then swiftly quenched star formation, on average. Thus, high SFR in high $\sigma$ ETGs plus stronger quenching winds lead to less chemical evolution and slightly lower [Fe/H] compared to low $\sigma$ galaxies, yielding a slightly negative trend in Fig. 9. These high $\sigma$ galaxies have higher light element abundances (Fig. 10), but one could posit that the time delay between Type II supernova enrichment and delayed Type Ia supernova enrichment makes enriched gas produced by the latter more susceptible to ejection by galactic winds.
Figure 23: Comparison of our results with previous work for log $\sigma$ - log age (top panel) and log $\sigma$ - [Fe/H] (bottom panel) relations. Note that the age plotted here for this work is the mean of age obtained from the 3rd annulus of each galaxy for each $\sigma$ bin.
@@
In Fig. 10 we see a trend of increasing abundance with $\sigma$ for all light elements explored in this work (C, N, Na, and Mg). Previous work on elemental abundances in ETGs shows a very similar trend (Parikh et al., 2021; Johansson et al., 2012; Conroy et al., 2014; Graves et al., 2007). Here we briefly discuss each element separately as each of them has different origin.
@@
Nitrogen primarily originates from the CNO cycle in stars, utilizing carbon, while carbon is predominantly produced via the triple alpha process, with supernovae contributing to both elements. Stellar winds in late evolutionary stages enrich the interstellar medium with these elements. Carbon production dominates in lower-mass stars (1–3 M), whereas nitrogen requires higher-mass stars (4–8 M) (Chiappini et al., 2003). In Milky Way stars, [C/Fe] remains near zero across [Fe/H], while [N/Fe] increases with metallicity (Jönsson et al., 2018). Our analysis reveals nitrogen under-abundance by approximately 0.1 dex relative to carbon and solar ratios, consistent with findings by Conroy et al. (2014) for carbon and nitrogen abundances and gradients.
@@
Not included in the models explicitly is the effect of first dredge-up on the first-ascent red giant branch. Nitrogen abundance increases by $\sim 0.4$ dex relative to C on the upper RGB [see, e.g., Fig 2 of Martig et al. (2016)]. If this effect were included, nitrogen’s spectral signatures would strengthen and the inferred N abundance would drop even lower in Fig. 11, putting [N/C] and [N/Fe] significantly low compared to the Milky Way. The effect is included implicitly, in that Milky Way stars were used to empirically zero the spectral library, and this effect is built in, masking as a surface gravity effect. That said, and keeping also in mind the N-rich globular cluster system in M31 that contrasts so strongly with our own cluster system (Worthey, 1998) and the small dynamic range of current galaxy formation model predictions (Fig. 33), it is safe to say that N is telling us much about chemical evolution that we are not yet equipped to interpret.
We see very similar abundances trajectories for sodium and magnesium (see Fig. 10). Although these two elements have different nucleosynthetic origins, both Na and Mg are mostly sourced from type II supernova (SNII) explosions (Marigo, 2001; Trager et al., 2000a) with significant contributions from the $\alpha$-rich freezout phase of the explosion (Woosley & Weaver, 1995), so their nucleosynthetic origin site is similar. The ‘locked-in’ state of ETGs in terms of chemical abundances preserve this signature of Na and Mg over time.
@@
A remarkable finding from this work is shown in Fig. 11. The effect shows more clearly because we did not stack spectra. What we see is that there is a very prominent increase in slope of chemical abundances when plotted against $\sigma$ for log($\sigma$)>>2.0. This slope change seems not to have been mentioned in previous literature on this subject. Possibly related, however, is that Gallazzi et al. (2005) found a transition from low mass to massive galaxies in the mass-metallicity relation at around log($\sigma$)=2.3. Evidently, low mass galaxies are less efficient in producing light metals and retaining them compared to massive galaxies. Different factors could lead to this, including star formation efficiency, galactic winds, and the IMF.
Figs. 12 (grouped by annulus) and 13 (third annulus per galaxy) depict elemental abundance dependence on [Fe/H]. A slight negative trend is observed, with light metal abundances decreasing as [Fe/H] increases. This trend in ETGs mirrors the Milky Way’s for magnesium but differs qualitatively for [C/R], [N/R], and [Na/R]. Simulations by Timmes et al. (1995) indicate that carbon and nitrogen remain constant over a wide [Fe/H] range, while sodium and magnesium decrease with increasing [Fe/H]. The decreasing trend of light metal abundances with [Fe/H] likely stems from iron production in type Ia supernova explosions (SNIa), which reduces [X/Fe] while increasing [Fe/H]. For magnesium, Sybilska et al. (2018) and Conroy & van Dokkum (2012) report a consistent decrease in [Mg/Fe] with increasing [Fe/H], corroborated by Alibés et al. (2001).
These chemical signatures of increasing light element abundance and decreasing (or flat) [Fe/H] with $\sigma$, coupled with high SFR and strong outflow, are consistent with ‘inside-out’ growth scenario during their formation. Their central regions likely underwent a brief, intense burst of star formation, rapidly enriching them in light metals and quenching soon after, leading to the high [X/Fe] ratios. More information can be obtained by studying the radial gradients in the abundances of iron-peak elements and other light metals.
5.2 Radial Gradients in Age and Abundances
@@
The radial gradients in different stellar population parameters provide insight into the formation and evolution of galaxies. In this work, in addition to looking at the trends of stellar population parameters with $\sigma$, we also investigated their radial gradients trend with $\sigma$.
@@
In Fig. 14 we show how the mean age for each $\sigma$ bin changes radially from the center of the galaxy to 0.65 Re. We find a modest radial gradient of $\sim$-0.04 log(Gyr)/log(R/Re) in mean age across different $\sigma$ bins, but the scatter is much larger than the mean trend. This result is also validated through Fig. 15. Mehlert et al. (2003) finds a median value of 0.0 in radial gradient in age for their Coma early-type galaxies. Sánchez-Blázquez et al. (2007) reports that 10 out of their 11 galaxies showed zero radial gradient in age. On the top panel of Fig. 24, we have compared our log $\sigma$ - $\nabla_{r}$ log age relation with some of the previous work by Kuntschner et al. (2010), Rawle et al. (2008), and Parikh et al. (2021). Our result, showing a nearly flat radial gradient in age, is consistent with these results within the scope of error. Most of the previous studies on radial gradient in age consider ages out to R=Re, but in this work we calculated spatial information to R=0.7 Re. A lack of age gradient might imply that stellar populations in ETGs were formed in bursts that spanned large swaths of galactocentric radius. However, it needs to be remembered that our results apply to the average. Few galaxies conform to the average. Slightly more than half are younger in the outskirts (0.7 Re for this study) and slightly less than half are younger in the center. Only a minority of galactic accretion events drive gas to galaxy centers, it seems. This result of nearly flat radial gradient in age (on average) indicates that stars likely formed in the inner regions and were subsequently mixed outward via dynamic processes, such as mergers and radial migration. These rapid, violent events in massive galaxies likely homogenized stellar populations across their radii, diminishing initial age gradients.
@@
We find radial gradients in [Fe/H] $\approx$-0.12 [dex per decade, [Fe/H]/log(R/Re)] for the middle $\sigma$ bin with substantial dependence on $\sigma$ (see Fig. 17). Mehlert et al. (2003) reports a value of -0.16 for radial gradient in [Z/H] (their Figure 7) and Parikh et al. (2021) shows an average radial gradient of around -0.19 in [Z/H] (their Figure 13). An unambiguous comparison was not possible between our work and theirs as we have used [Fe/H] as proxy for metallicity and most of the other work use [Z/H] as proxy for metallicity. However, direct comparison with results from Kobayashi (2004) are shown in Fig. 24. The bottom panel of Fig. 24 shows that our radial gradient in [Fe/H] is on average a factor of two shallower than Kobayashi (2004). Earlier estimates averaged a slope of 0.2-0.2 dex per decade for medium and large ETGS (Henry & Worthey, 1999). We see a considerable scatter around zero (see top panel of Fig. 17) with values as negative as -0.8 [Fe/H]/log(R/Re) and as positive as 0.4 [Fe/H]/log(R/Re) (not considering outliers). Our trend of radial [Fe/H] gradient with $\sigma$ qualitatively matches with Spolaor et al. (2009) at low $\sigma$ values (log($\sigma$)<<2.0) and with Bekki & Shioya (1999) for higher $\sigma$ values. A negative [Fe/H] gradient favors an ‘inside-out’ mode of chemical evolution in ETGs if star formation and enrichment occur on a timescale similar to the accretion itself. This conclusion is similar to that of Goddard et al. (2017) where they find similar results with 505 ETGs. The [Fe/H] radial gradient reported here could also be explained as a consequence of hierarchical merging events. If they involve gas, mergers can trigger intense central starbursts, which significantly enrich the core regions with metals leading to a metallicity distribution where the central areas exhibit higher metal content than the outer regions (Kobayashi, 2004). If not subsumed into star formation right away, gas loses energy and angular momentum, funneling into the central regions where concentrated star formation enriches the core with metals more significantly than the outskirts (Montes et al., 2014). The gradual increase of [Fe/H] radial gradient with $\sigma$ can also be explained by the amplification of these processes for larger galaxies. The monolithic collapse hypothesis, on the other hand, predicts a much higher [Fe/H] gradient (Kawata, 2001; Pipino et al., 2010) than what is reported here. If monolithic collapses are a common formation channel, then we also require major “dry” merger events to explain the much flatter slope in [Fe/H].
Figure 24: Comparison of our results with various other previous work for log $\sigma$ - $\nabla_{r}$log(age) (top panel) and log $\sigma$ - $\nabla_{r}$[Fe/H] (bottom panel) relations.
@@
Using a linear scale for R/Re, the negative radial gradient in age is 0.04 dex per R/Re for the lowest $\sigma$ bin and remains constant at an increased value of 0.11 dex per R/Re for the rest of the $\sigma$ bins. We quantify the goodness of fit for both logarithmic and linear cases using the coefficient of determination (the R2 value). The R2 values for each $\sigma$ bin for the linear case are (in order of increasing $\sigma$): 0.52, 0.93, 0.97, 0.94, and 0.95. On the other hand, for the logarithmic case, the R2 values are: 0.26, 0.70, 0.77, 0.71, 0.85. Thus, in general, the linear case shows better fit. Similar analyses were also performed for the radial gradients in [Fe/H]. The negative radial gradient in [Fe/H] (using linear scale for R/Re) increases from 0.12 dex per R/Re to 0.23 dex per R/Re with R2 values of 0.96, 0.97, 0.97, 0.98, and 0.98 for different $\sigma$ bins (in order of increasing $\sigma$). Using the logarithmic scale the R2 values are 0.79, 0.81, 0.91, 0.93, and 0.84, again showing better fit for the linear case. Thus for both age and [Fe/H] we find that a linear physical scale in kpc connects to the spatial dependencies of chemical evolution better than a logarithmic one. We speculate that power-law dependencies such as depth of potential well or dynamical crossing time are less important than turbulence or other processes with a relatively fixed sphere of influence.
@@
Fig. 18 shows the radial trends of all of the light elements explored in this work. We discuss each of the elements individually.